### v2

# What Changed and What Could Have Changed? State-Change Counterfactuals for Procedure-Aware Video Representation Learning

Abstract
Understanding a procedural activity requires modeling both how action steps transform the scene, and how evolving scene transformations can influence the sequence of action steps, even those that are accidental or erroneous. Existing work has studied procedure-aware video representations by proposing novel approaches such as modeling the temporal order of actions and has not explicitly learned the state changes (scene transformations). In this work, we study procedure-aware video representation learning by incorporating state-change descriptions generated by Large Language Models (LLMs) as supervision signals for video encoders. Moreover, we generate state-change counterfactuals that simulate hypothesized failure outcomes, allowing models to learn by imagining the unseen “What if” scenarios. This counterfactual reasoning facilitates the model’s ability to understand the cause and effect of each step in an activity. To verify the procedure awareness of our model, we conduct extensive experiments on procedure-aware tasks, including temporal action segmentation, error detection, action phase classification, frame retrieval, multi-instance retrieval, and action recognition. Our results demonstrate the effectiveness of the proposed state-change descriptions and their counterfactuals, and achieve significant improvements on multiple tasks. We will make our source code and data publicly available soon.
Figure 1: Illustration of “what change” (actual action-induced transformations) and “what could have changed” (hypothetical deviations) for both clip- and video-level procedures. (a) The clip narration “Scoop oatmeal with a spoon and drop it into a bowl of water” changes the scene’s Before and After states at the short-term clip-level. (b) Illustrates the hypothesized deviation, i.e., State-change Counterfactual, that could have happened and changed the scene differently at the clip level. (c) Illustrates the counterfactuals for video-level procedures, which depicts what could happen due to possible previous mistakes, for example, Missing-step Counterfactual and Misordered Counterfactual.
[…]
1 Introduction
Procedural activities, such as following a cooking recipe [gtea, breakfast, egoper, grauman2024ego] or assembling a piece of furniture [miech2019howto100m, sener2022assembly101], consist of sequences of interconnected steps that often happen in a specific and logical order to achieve a desired outcome. Understanding these activities from video data is essential for a variety of applications, including video retrieval [dong2019dual, miech2019howto100m, gabeur2020multi, Wray_2021_CVPR], intelligent collaborative agents [kung2023riskbench, wang2024mosaic, banerjee2024ask, ye2024morpheus, jenamani2024feel], and robot learning from human demonstration [7102751, shao2021concept2robot, xiao2022masked, zakka2022xirl, xu2023xskill, zhu2024orion, jain2024vid2robot, ren2025motion]. In contrast to general video action recognition that focuses on a single step in a short clip [carreira2017quo, feichtenhofer2019slowfast, fan2021multiscale, wang2021actionclip, tong2022videomae, Kung_2024_CVPR], procedure-aware video understanding requires capturing both the “what changed” (actual action-induced state transformations) [liang2022var, souvcek2023genhowto], as shown in Figure 1 (a), and “what could have changed” (hypothetical deviations) [HARRIS1996233, roese1997counterfactual, schulam2017reliable, zhang2021if], as shown in Figures 1 (b) and (c). These two capabilities play crucial roles in understanding long-form procedures where each action step can have a sequential effect on the others.
Procedure-aware video representation learning can enable various downstream long-form video tasks, such as temporal action segmentation [gtea, breakfast, sener2022assembly101, egoper] and error detection [sener2022assembly101, egoper]. There have been many approaches proposed to learn procedure-aware representations, including learning spatiotemporal features [carreira2017quo, feichtenhofer2019slowfast, fan2021multiscale, tong2022videomae], using action labels as supervision [miech2020end, lin2022learning, xiao2022hierarchical, zhong2023learning], incorporating temporal order of steps [jenni2020video, zhong2023learning], and consulting external activity procedure knowledge databases [zhou2023procedure] such as WikiHow [koupaee2018wikihow]. However, these approaches often struggle to explicitly model procedural knowledge that relies on how scene states evolve over time and how they could have evolved if the precondition or result of an action step had differed. For instance, the procedure of “Make Oatmeal” in Figure 1 (c) cannot be finished correctly if the state of the bowl remains “Full of oatmeal and without water” or “Full of water without oatmeal” because they do not meet the precondition of the later action step, “Heat with microwave”. Capturing these subtle transitions requires state-change reasoning to understand how an action transforms an environment or objects, and counterfactual analysis to foresee alternative outcomes for the same key action.
In this paper, we propose a hierarchical procedure-aware video representation learning framework that leverages state-change and their counterfactuals generated by a Large Language Model (LLM) [dubey2024llama]. We use the actual state-change descriptions to temporally contrast features, and the counterfactuals as the negative exemplars in contrastive learning [oord2018representation, chen2020simple, 10.5555/3495724.3497291] to enhance the modeling of clip-level and video-level procedure transformations. Specifically, at the clip level, we first capture before and after state that describe the change of environments or objects to identify local action-induced state changes. Moreover, we also incorporate state-change counterfactuals that estimate a hypothesized state if the action step were to fail for any reason. For example, if the action were “make coffee,” a possible state counterfactual could be “cup shards scattered on the floor.” We then use temporal contrastive learning to encourage the model to learn a representation space in which visual features of later frames in short clips and after states are close together, and in which the early frames, before state, and state counterfactual are far away, and vice versa. At the video level, we further incorporate counterfactual reasoning by generating missing-step and misordered, enabling a more comprehensive understanding of the entire procedure in a long-form video.
We evaluate the learned procedure-aware video representations in six key procedural and short-term video understanding tasks—error detection, temporal action segmentation, multi-instance retrieval, action recognition, action phase classification, and frame retrieval—and show that they significantly enhance performance compared to strong baselines. To validate our approach, we incorporate state changes and their counterfactuals into the learning process and pre-train our model on Ego4D [grauman2022ego4d, kevin2022egovlp], a large-scale daily activity dataset filtered to include clean and aligned short clips along with their corresponding clip-level narration text and video-level summary text. We then evaluate our model on various procedure-aware downstream tasks, including long-form video tasks like temporal action segmentation and error detection, and short-term action recognition and retrieval. In addition, we conduct a comprehensive ablation study to highlight the impact of learning state changes and their counterfactuals, demonstrating their importance in enhancing procedural video understanding. Our contributions are summarized as follows:
  1. 1.
    We introduce a novel approach to video representation learning that explicitly incorporates state changes and their counterfactuals to enhance procedure awareness.
  2. 2.
    We design a hierarchical learning framework that effectively integrates state-change and counterfactual descriptions into frame-, clip-, and video-level feature alignment.
  3. 3.
    We demonstrate the effectiveness of our learning approach with state-of-the-art performance on several procedure-aware video downstream tasks and offer comprehensive analysis and discussion.
2 Related Work
2.1 Video Procedure-Aware Representations
Procedure understanding, which involves recognizing and localizing individual action steps and reasoning about their causal relationships, is a fundamental challenge in computer vision, and is important for applications such as step recognition [tang2019coin, damen2020epic, bansal2022my], temporal action segmentation [gtea, breakfast, sener2022assembly101, chen2025atars], and error detection [egoper]. Early work proposed learning procedure-aware video representations via exploiting the temporal consistency within videos [wang2015unsupervised, bansal2022my, xiao2022hierarchical, zhang2022unsupervised, qing2022learning] and temporal order [jenni2020video] as pre-text tasks. Some work incorporates task graphs as guidance to learn procedure-aware representations, either by constructing graphs based on knowledge bases [zhou2023procedure] or through learning [narasimhan2023learning, mavroudi2023learning, bansal2024united, seminara2024differentiable, ashutosh2024video]. Recently, several papers have attempted to align video features in datasets such as HowTo100M [miech2019howto100m] or Ego4D [grauman2022ego4d] with text descriptions extracted through automatic speech recognition [miech2020end, mavroudi2023learning], refined subtitles [lin2022learning] with the external database WikiHow [koupaee2018wikihow, mavroudi2023learning], or manually labeled annotations [kevin2022egovlp, zhong2023learning, ashutosh2023hiervl]. These text descriptions offer rich context of action steps and temporal relations, opening the door for procedure awareness learning. For example, Br-prompt [li2022bridge] generates text prompts from ground truth action labels to learn ordinal temporal features. PVRL [zhong2023learning] formulates future action steps as next-token prediction [devlin2019bert], enhancing temporal order between actions. HierVL [ashutosh2023hiervl] proposes hierarchical video-text alignment for both short clips and long videos.
However, these papers fall short of explicitly learning state changes—the causal relationships between action and state—which is essential to procedure understanding [souvcek2023genhowto, niu2024schema]. Moreover, they overfit the seen correctly executed actions, which hinders the understanding of activity procedures and the ability to recognize the deviation of actions, such as erroneous action steps. In this work, we propose to incorporate state-change descriptions and counterfactuals to force our model to learn about action transformations and hypothetical outcomes that facilitate “learning from hypothetical failures” for procedure awareness.
Figure 2: Illustration of our learning framework. (a) We train frame features by temporally contrasting neighboring and distant frames by incorporating Before, After, and State-change Counterfactuals (SC-CF). (b) We align clip features with narration features. (c) Clip features are aggregated with the aggregator (Agg) as video features. Then the video features are contrasted positively with summaries and negatively with Missing-step Counterfactuals (K-CF) and Misordered Counterfactuals (M-CF). Note that all text features are extracted with the frozen text encoder.
2.2 State-Change Understanding
The idea of visual state changes has been explored extensively in the computer vision community, sometimes using different terms such as transformations between states [wang2016actions, liang2022var] or state-modifying actions [souvcek2022multi, souvcek2022look, niu2024schema, soucek2024multitask]. For example, papers have studied object state change localization [liu2017jointly, alayrac2017joint, souvcek2022look, saini2023chop, xue2024learning] and generation [soucek2024multitask], and action recognition with state changes [fathi2013modeling, wang2016actions, saini2023chop]. Procedure planning, which aims to generate effective plans for task completion based on start and goal visual states, is another task that heavily incorporates the concept of understanding state changes [chang2020procedure, lu2022neuro, wang2023event, li2023skip, niu2024schema, nagasinghe2024not]. We were also inspired by SCHEMA [niu2024schema] to use LLMs to generate before- and after-state descriptions of actions. However, three key factors sets our work apart from the existing work. First, unlike existing work that mainly focuses on a specific task such as procedure planning [niu2024schema], we aim to design a more general video representation learning framework for various downstream tasks. Second, prior work does not incorporate hypothetical outcomes, i.e., counterfactuals, which are essential for procedure causal reasoning. Last but not least, prior work often only studies interacted objects’ state changes. In contrast, our generated descriptions can cover state changes of objects, humans, and the surrounding environment. We provide extensive examples in our supplementary materials.
2.3 Video Representations with Counterfactuals
Counterfactual example reasoning that helps models understand cause-and-effect and predict outcomes [scholkopf2021toward, yang2021causalvae, zang2023discovering] has demonstrated strong generalization and robustness on various tasks including video question-answering [CLEVRER, li2022representation, zhang2024if] and compositional reasoning [sun2021counterfactual, lai2024improving]. Extensive research has proposed to enhance the causality in visual learning by creating counterfactual examples with image data augmentation [zhong2020random], text counterfactuals based on rule-based [shi2018learning], random word substitution [roth2023waffling], language models [doveh2023teaching, wang2023vilta, xie2023ra, zhang2024countercurate], and synthetic image-text pairs generated by a simulator [cascante2023going] or generative model [shi2023logoprompt, le2023coco, lai2024improving], such as stable diffusion [rombach2022high]. However, existing work mainly focuses on creating counterfactuals for images or short clips and does not consider the procedural nature of activities. In this work, we propose state-change counterfactuals at both the clip and video levels, highlighting the action transformations and procedure evolution.
3 Proposed Method
We first present the text generation process for state changes and state-change counterfactuals. Then, we provide an overview of the clip-video hierarchical representation learning framework. Finally, we present our pretraining optimization that incorporates the generated texts to learn procedure-aware representations.
3.1 State Change & Counterfactual Generation
We use the LLM Llama 3 [dubey2024llama] to generate clip-level state descriptions, i.e., before state, after state and state-change counterfactual, and to generate video-level summary counterfactuals, i.e., missing step and misordered steps. In the following, we introduce the prompts for text generation.
Clip-Level State Changes and their Counterfactuals. We feed the narration annotated for each short clip in Ego4D [grauman2022ego4d, kevin2022egovlp] into Llama and ask it to generate a state-change description based on the given narration. Specifically,
    Generate [Before] describing the    scene before the action is    performed and [After] describing    the scene changed by the action.For example, given the narration “C picks a bag of clothes from the floor,” where C is the camera-wearer, Llama might generate [Before]: “The floor is cluttered with clothes” and [After]: “The bag of clothes is now in C’s hand, with the surrounding area slightly rearranged”. Then, we generate multiple state-change counterfactuals (SC-CF) using another prompt with the same action context,
    Now, based on the state changes you    generated, create 3 state-change    counterfactuals [SC-CF] depicting    scenes with incomplete or    incorrectly completed actions.The generated results [SC-CF] might be “Clothes remain scattered on the floor” or “A small pile of clothes sits amidst remaining clutter”.
Video-Level Counterfactual Generation. We feed all clip narrations in a long video and their video-level summaries annotated in Ego4D [grauman2022ego4d, kevin2022egovlp] to Llama. We ask the LLM to generate summary counterfactuals for missing key steps [K-CF] in which some crucial action steps (clip narrations) are missing,
    Generate 10 distinct    counterfactuals [K-CF] by    taking out some key actions.and misordered [M-CF] counterfactuals in which the order of key action steps is permuted,
    Generate 10 distinct    counterfactuals [M-CF] by    perturbing the order of actions.
In this way, we generate diverse counterfactuals that can facilitate models to learn from hypothetical and unseen scenarios. As an example, consider the video summary “Make omelet with rice” in Figure 2. The procedure consists of a sequence of actions, [“Crack eggs,” “Whisk eggs,” “Fry eggs,” “Fry rice”, …]. The video-level state-change counterfactuals could be missing action step “Whisk eggs”, which might lead to state “Clumpy eggs”, while if one misordered action step “Fry rice” first and then move to other steps, e.g., “Crack eggs”, “Whisk eggs”, and so on, it could lead to state “Burnt rice”. We provide more examples of the generated descriptions in the supplementary materials.
3.2 Preliminary: Hierarchical Video
Feature Learning
In this work, we adopt HierVL [ashutosh2023hiervl], a framework that learns hierarchical video-language representations at two temporal scales, clip-level and video-level. The clip-level narrations and video-level summaries are annotated by Ego4D/Egoclip [grauman2022ego4d, kevin2022egovlp]. Note that since HierVL trains both a visual and text encoder, it requires vision-to-text and text-to-vision alignment. Below we describe the former, and the latter is defined symmetrically. Formally, let a video be divided into $M$ short clips {$v_{1},\ldots,v_{M}$} with corresponding narration texts {$t_{1},\ldots,t_{M}$}, and let $S$ be the video-level summary for the entire video. HierVL’s vision-to-text training objective consists of two parts: clip-level alignment and video-level alignment losses.
Clip-Level Alignment (Child Loss) is a contrastive loss that encourages matched $(v_{i},t_{i})$ pairs to have high similarity relative to mismatched pairs. Specifically, we use a softmax-based contrastive objective,
$\mathcal{L}_{{v2t}}=\frac{-1}{|B|}\sum_{i\in B}\log\frac{\sum_{p\in P(i)}\exp(v_{i}^{T}t_{p}/\tau)}{\sum_{j\in B}\exp(v_{i}^{T}t_{j}/\tau)},$ (1)
where $B$ is the mini-batch, $v_{i}$ and $t_{j}$ are visual and textual embeddings, respectively, and $P(i)$ denotes the positive samples of the $i^{\text{th}}$ video clip. HierVL [ashutosh2023hiervl] leverages a variant of the action-aware loss, EgoNCE [kevin2022egovlp], in order to mine hard positives and negatives; please refer to [ashutosh2023hiervl, kevin2022egovlp] for more details about this sampling. The overall loss at the child level is thus $\mathcal{L}_{\text{child}}=\mathcal{L}_{{v2t}}+\mathcal{L}_{{t2v}}$, where the latter is defined symmetrically.
Video-Level Alignment (Parent Loss) captures long-range context. Clip features ${v_{1},...,v_{m}}$ and text features ${t_{1},...,t_{m}}$ are aggregated into single video-level representations $V_{i}$ and $T_{i}$, respectively,
$V_{i}$ $=\text{Agg}({v_{1},\ldots,v_{m}}),$
$T_{i}$ $=\text{Agg}({t_{1},\ldots,t_{m}}),$
where $\text{Agg}(\cdot)$ is an aggregator function implemented with a self-attention transformer. This parent loss seeks to align the video-level features to the text summary provided by Ego4D,
$\mathcal{L}_{\text{parent}}^{SV}=\frac{-1}{|B|}\sum_{i\in B}\log\frac{\sum_{p\in P(i)}\exp(V_{i}^{T}S_{p}/\tau)}{\sum_{j\in B}\exp(V_{i}^{T}S_{j}/\tau)},$ (2)
where the superscript $SV$ emphasizes that this is the vision-to-text alignment loss, and $\mathcal{L}_{\text{parent}}^{ST}$ is defined analogously for text-to-text alignment. The overall loss at the parent level is thus $\mathcal{L}_{\text{parent}}=\mathcal{L}_{\text{parent}}^{SV}+\mathcal{L}_{\text{parent}}^{ST}$.
Table 1: Temporal action segmentation results on the GTEA and EgoPER datasets. The underline denotes the second best. Text encoder denotes the VLM model with a trainable text encoder. Following prior work [10377231] we also present the average across all metrics (Avg).
GTEA EgoPER
Method Pretraining Data F1@10 F1@25 F1@50 Edit Acc Avg F1@50 Edit Acc Avg
Br-prompt [li2022bridge] GTEA 94.1 92.0 83.0 91.6 81.2 88.4 - - - -
I3D [carreira2017quo] Kinetics 90.1 88.8 79.2 84.6 79.7 84.5 48.8 71.9 73.9 64.9
CLIP [radford2021learning] WIT [srinivasan2021wit]+Text encoder 88.5 86.2 77.6 87.1 75.6 83.0 44.2 71.2 70.8 62.1
MIL-NCE [miech2020end] HowTo100M 67.9 61.3 44.6 67.9 58.3 60.0 47.3 69.1 73.6 63.3
PVRL [zhong2023learning] HowTo100M 85.2 82.6 72.2 81.1 71.2 78.5 45.6 73.2 73.4 64.1
HierVL [ashutosh2023hiervl] Ego4D+Text encoder 90.4 88.5 81.2 86.7 78.5 85.1 52.6 73.0 77.3 67.6
Ours Ego4D 89.8 89.1 81.6 86.8 80.0 85.5 54.4 74.1 79.0 69.2
3.3 Pretraining Objective:
State Change & Counterfactual
We use a hierarchical contrastive pre-training framework consisting of three timescales: frame-level, clip-level, and video-level.
Frame-Level Alignment. At the frame level, the model is supervised by our proposed before-state loss $\mathcal{L}_{\text{before}}$ and after-state loss $\mathcal{L}_{\text{after}}$, enhancing action-induced transformation for procedural understanding. The mathematical formulation of each is [10.5555/3495724.3497291],
$\mathcal{L}$ $=\frac{1}{|B|}\sum_{i\in B}\frac{-1}{|P(i)|}\sum_{p\in P(i)}\log\frac{\exp(f_{i}^{T}z_{p}/\tau)}{\sum_{n\in N(i)}\exp(f_{i}^{T}z_{n}/\tau)},$ (3)
where $B$ is the batch size, $f_{i}$ is the visual embedding of the $i^{\text{th}}$ frame, $z_{j}$ is either a visual or text embedding, $\tau$ is a temperature hyperparameter, and $P(i)$ and $N(i)$ denote the positive and negative samples of the $i^{\text{th}}$ frame, respectively. Given a set of $K=4$ frames sub-sampled from a video-clip, $L_{\text{before}}$ aims to align earlier-in-time frames along with the before-state text embeddings, while pushing them apart from later-in-time frames, the after-state, and counterfactual text embeddings. That is, $P(0)=\{f_{1},t_{b}\}$ and $N(0)=\{f_{3},t_{a},t_{cf}\}$, where $t_{b}$, $t_{a},$ and $t_{cf}$ denote the text embeddings corresponding to the before state, after state, and the counterfactuals, respectively. On the other hand, $L_{\text{after}}$ aims to align later-in-time frames to the after state while separating them from earlier-in-time frames, the before-state, and counterfactual text embeddings. In other words, $P(3)=\{f_{2},t_{a}\}$ and $N(3)=\{f_{0},t_{b},t_{cf}\}$. Additionally, other samples from the same mini-batch are included as negatives in the denominator of Eq. (3) in both losses, but we omit it in the notation above for simplicity.
Clip-Level Alignment. At the clip-level, we leverage the $L_{v2t}$ loss described in Section 3.2, which seeks to align the video-clip embeddings to their corresponding text narrations from EgoClip [kevin2022egovlp]. Note that since we do not train a text encoder, the symmetric $L_{t2v}$ loss is neglected here. For more details on this loss, see Section 3.2 and [ashutosh2023hiervl, kevin2022egovlp]. The resulting loss for the first two scales is therefore
$\mathcal{L}_{\text{child}}=\mathcal{L}_{v2t}+\lambda(\mathcal{L}_{\text{before}}+\mathcal{L}_{\text{after}}),$ (4)
where $\lambda$ is a hyperparameter controlling the strength of the state-change aware supervision.
Video-Level Alignment. The goal of this loss is to align video-level visual embeddings to summary text embeddings and to enhance procedural awareness using video-level counterfactuals. We first obtain video-level visual embeddings from clip-level embeddings by leveraging a self-attention block as described in Section 3.2. Then, using contrastive learning, each visual embedding is aligned to its corresponding summary text embedding and contrasted against text embeddings of the generated counterfactual K-CF and M-CF. Note that this allows us to decouple the summary and counterfactual losses; however, we find no noticeable performance difference when doing so and thus opt to use a joint loss in practice. The formulation is defined as,
$\mathcal{L}_{\text{parent}}$ $=-\sum_{i\in B}\log\frac{\sum_{p\in P(i)}\exp(V_{i}^{T}S_{p})}{\sum_{n\in N(i)}\exp(V_{i}^{T}S_{n})+\exp(V_{i}^{T}S_{n,w}^{cf})},$ (5)
where $V_{i}$ is the aggregated video-level visual embedding, $S_{j}$ is a summary text embedding and $S_{j,w}^{cf}$ are Misordered and Missing-step counterfactual text embeddings where $w\in\{1,\ldots,W\}$ is the total number of counterfactuals used, and $P(i)$ and $N(i)$ denote the positive and negative samples of the $i^{\text{th}}$ video, respectively. We also use a temperature hyperparameter $\tau$ as in Eq. (3), but we omit it in the equation above for simplicity of notation. Note that consistent with [ashutosh2023hiervl], in this case, the summation over positive samples is located inside the logarithm since the sampling strategy of [kevin2022egovlp] is used. However, prior work [10.5555/3495724.3497291] finds the summation outside the logarithm to be more effective and thus this is the de-facto choice of the frame-level loss in Eq. (3).
4 Experiments
4.1 Implementation Details
Pretraining. We pretrain our model on EgoClip [kevin2022egovlp], a dataset derived from Ego4D [grauman2022ego4d], with clean clip-level narrations and the video-level summaries. We follow [kevin2022egovlp, ashutosh2023hiervl] to use the same training split that has 3.8M clip narrations and 120K long-term video summaries for pretraining. We pretrain our model on 8 NVIDIA L40S GPUs, with batch size 18 on each GPU for 7 epochs. We follow HierVL [ashutosh2023hiervl] and train models on clip- and video-level alignment alternatively. Specifically, we perform video-level alignment as in Eq. (5) after every 5 mini-batches of clip-level alignment as in Eq. (4). The frame-level loss is computed during the clip-level training iteration. We set the learning rate and weight decay to 1e-5 and 1e-4, respectively. Note that each short clip is sub-sampled to 4 frames during training [kevin2022egovlp, ashutosh2023hiervl].
Text Generation. To generate all state-change descriptions and state-change counterfactuals, we use Llama 3.2 8B [dubey2024llama] for efficiency. We incorporate off-the-shelf text-encoder FLAVA [singh2022flava] to pre-extract features of the annotated narrations and summaries and our generated text descriptions.
Temporal Action Segmentation. We evaluate representations on temporal action segmentation, where models output the frame-wise action label for untrimmed, long videos, requiring the ability to reason about procedures and recognize the start and end of state transformations. We evaluate various representations by adopting the popular temporal action segmentation model ASFormer [chinayi_ASformer] that takes video features as input. We follow prior work [li2022bridge] for the same training strategy and evaluation protocol and metrics including F1 scores, Edit distance, and frame-wise Accuracy, on the cooking procedure datasets, Georgia Tech Egocentric Activities (GTEA) [gtea] and EgoPER [egoper] Dataset. GTEA consists of 28 egocentric instructional videos capturing daily kitchen activities at 15 frames per second. EgoPER contains 368 videos spanning 28 hours of cooking scenarios.
Error Detection. We evaluate representations on unsupervised error detection on the EgoPER dataset [egoper] where models are trained with error-free videos and output labels for each frame on whether it is an error during testing time. The error categories include step omission, step addition, step modification, step slip, and step correction. We follow [egoper] and use its EgoPED model [egoper], and equip it with different video representations. Specifically, the training process of EgoPED consists of two stages: action segmentation and prototype learning. In stage 1, the temporal action segmentation model, ASFormer, is trained to predict frame-wise action labels on the error-free set by taking pre-extracted video features as input. In stage 2, the model learns prototype features for each class of action step, e.g., “Take bowl from microwave,”, with contrastive step prototype learning [oord2018representation, egoper]. For each class of activity, multiple prototypes are extracted for error detection at inference. Note that each activity is trained independently, such as Make Coffee. During the inference, EgoPED calculates the similarity score of observed frame features and any learned prototypes. A threshold is further used to determine whether the observed test frame is erroneous.
Action Retrieval & Recognition. To demonstrate the applicability of the learned representations in the core video understanding tasks, we evaluate zero-shot multi-instance retrieval on Epic-Kitchens (EK) [Damen2018EPICKITCHENS] and zero-shot action recognition on Charades-Ego (CE) [sigurdsson2018charadesegolargescaledatasetpaired]. Following prior work [kevin2022egovlp, ashutosh2023hiervl], both tasks are implemented via text-to-vision alignment by selecting the textual label with the highest similarity score to the given visual sample.
Table 2: Error detection results on the EgoPER dataset. The underline denotes the second best. HTM denotes the HowTo100M dataset [miech2019howto100m]. Text denotes the VLM model with a trainable text encoder.
Quesadilla Oatmeal Pinwheel Coffee Tea All
Method Pretraining Data EDA AUC EDA AUC EDA AUC EDA AUC EDA AUC EDA AUC
Random - 19.9 50.0 11.8 50.0 15.7 50.0 8.20 50.0 17.0 50.0 14.5 50.0
HF2-VAD [liu2021hybrid] - 34.5 62.6 25.4 62.3 29.1 52.7 10.0 59.6 36.6 62.1 27.1 59.9
S3R [wu2022self] Kinetics+I3D 52.6 51.8 47.8 61.6 50.5 52.4 16.3 51.0 47.8 57.9 43.0 54.9
I3D [carreira2017quo] Kinetics 62.7 65.6 51.4 65.1 59.6 55.0 55.3 58.3 56.0 66.0 57.0 62.0
CLIP [radford2021learning] WIT [srinivasan2021wit]+Text 77.6 67.2 69.6 67.5 66.9 59.5 68.5 69.0 75.6 57.7 71.6 64.2
MIL-NCE [miech2020end] HTM 77.3 59.8 69.8 61.5 65.7 53.0 68.0 67.9 69.8 61.5 70.1 60.7
PVRL [zhong2023learning] HTM 75.7 70.0 71.2 53.5 65.5 65.5 67.5 65.4 76.4 65.2 71.3 63.9
HierVL [ashutosh2023hiervl] Ego4D+Text 77.9 70.2 70.8 66.4 65.2 58.1 67.4 69.8 75.1 66.4 71.3 66.2
Ours Ego4D 78.9 63.7 71.6 46.1 68.3 53.8 68.3 68.0 76.6 70.9 72.7 60.5
Action Phase Classification & Frame Retrieval. To assess more fine-grained and short-term procedure awareness, we test representations on action phase classification and zero-shot frame retrieval on the Align-Ego-Exo (AE2) dataset [xue2023learning]. The AE2 dataset contains egocentric and exocentric videos displaying the actions Break Eggs, Pour Milk, Pour Liquid, and Tennis Forehand. Each action is divided into two to four phases [xue2023learning]. We follow [xue2023learning] and train an SVM to predict the per-frame action phase labels for the classification task and use nearest neighbors for the zero-shot retrieval task. Note that we merge the validation and test sets for more robust results.
Baselines. We evaluate the popular I3D and CLIP feature [carreira2017quo, radford2021learning] commonly used in long-form video tasks and procedure-aware representations with publicly available pretrained model weights, including MIL-NCE [miech2020end], Bridge-prompt [li2022bridge], and PVRL [zhong2023learning]. In addition, we evaluate the VLM HierVL [ashutosh2023hiervl] with only its video encoder.
4.2 Temporal Action Segmentation Results
Table 1 presents results on temporal action segmentation. The first line of the table shows Br-prompt [li2022bridge] which is pre-trained on the target dataset (GTEA) and thus serves as an upper bound on performance. In our results, we found that MIL-NCE performs significantly worse than others, especially on GTEA. We assume this is because MIL-NCE is trained with low-quality automatic-speech-recognition (ASR) text and imprecise alignment between video and ASR sentences, unlike PVRL [ashutosh2023hiervl] which leverages pseudo-labels generated by CLIP [radford2021learning]. We further compare our model with the VLM HierVL. Even though HierVL is trained with both video and text encoders, which require considerably more computational resources, our model outperforms HierVL in most metrics on both datasets. Furthermore, the performance gap against all non-VLM models is larger. This highlights the effectiveness of our proposed state-change descriptions and their counterfactuals on long-term procedure understanding.
4.3 Error Detection Results
Table 2 presents results of the error detection benchmark on EgoPER [egoper]. We first include results reported in EgoPER as a reference in the top group (rows 1-3) instead of a direct comparison. In the remaining groups (rows 4-9), we report the performance of representation learning approaches, tested using the same downstream architecture [chinayi_ASformer]. We observe that all procedure-aware representations outperform general visual representations, such as I3D and CLIP, highlighting the importance of procedure awareness in the context of error detection. Furthermore, our proposed method surpasses the state-of-the-art in EDA [egoper], even outperforming the VLM HierVL.
Moreover, we achieve competitive performance on the AUC metric. Note that although we follow prior work [egoper] to include frame-wise metric AUC, this metric may not be effective in measuring the performance and the procedure awareness of representations. As discussed in EgoPER [egoper], the heuristic random method can achieve competitive results in AUC but fail to localize erroneous segments as shown in the major metric EDA. This is also true when comparing I3D and earlier methods, such as HF2-VAD [liu2021hybrid] and S3R [wu2022self]. We hope our extensive experimental results can inspire future work to study more effective evaluation metrics for error detection. In summary, our representation learning method achieves state-of-the-art performance in EDA and competitive performance in AUC, highlighting the strong effectiveness of counterfactual reasoning in procedure awareness.
Table 3: Multi-Instance Retrieval (MIR) on Epic-Kitchens (EK) and Zero-shot Action Recognition on Charades-Ego (CE). We report PT ckpt mAP for the Action Recognition task (see [ashutosh2023hiervl]).
MIR (EK) Action Rec. (CE)
Method mAP nDCG mAP
CLIP [radford2021learning] 8.4 15.3 20.5
MIL-NCE [miech2020end] 5.8 10.3 6.9
PVRL [zhong2023learning] 6.1 11.4 7.9
Ours 15.7 22.6 24.8
4.4 Action Retrieval & Recognition Results
Table 3 presents benchmark results on the action retrieval and recognition. We found existing procedure-aware representations perform significantly worse on short-term action recognition tasks. Yet, our representations trained with both short- and long-term state-changes and counterfactuals demonstrate strong applicability to the popular video recognition tasks and datasets.
4.5 Action Phase Classification & Retrieval Results
Table 4 shows that our model outperforms others on average across actions on both the ego+exo and ego settings, demonstrating its effectiveness in short-term procedures and strong view-point generalization. Interestingly, CLIP [radford2021learning] demonstrates strong performance. We conjecture this is due to its focus on short-term Full results with individual actions are available in Section LABEL:subsec:full in the supplementary material.
Table 4: Action Phase Classification and zero-shot Frame Retrieval results on the Align-Ego-Exo dataset [xue2023learning].
Classification (F1) Retrieval (mAP)
Method ego+exo ego ego+exo ego
CLIP [radford2021learning] 59.5 61.6 64.4 68.0
MIL-NCE [miech2020end] 53.0 54.2 59.5 63.0
PVRL [zhong2023learning] 59.4 62.7 61.6 66.6
Ours 61.3 64.8 64.9 70.3
[…]
ID Before & After SC-CF K-CF M-CF GTEA (AS) EgoPER (ED)
[…]
4.6 Ablation Study
In Table 5, we study the effectiveness of state changes (Before and After, state-change counterfactuals (SC-CF), and video-level state-change counterfactuals (Missing-step CF and Misordered-step CF) by ablating each of them and evaluating on temporal action segmentation with GTEA and error detection with EgoPER. Note that ID 1 can be seen as Weak HierVL as it is trained with only the video encoder. In addition, since Weak HierVL only uses narrations from other videos in a batch as negatives, it can be seen as a random description approach. All ablations (including Ours) are pre-trained with a batch size of 12 on each GPU.
Temporal Action Segmentation. One interesting insight on GTEA is that using only clip-level state changes and their counterfactuals (ID 3) achieves competitive performance on F1 and Edit metrics, which outperforms most other ablated models. Also, using only one type of video-level counterfactual (ID 4 and 5) may mislead the model to overfit a certain type of counterfactual, resulting in lower performance. By considering all the proposed states and counterfactuals together, we show that they are complementary to each other, thus achieving the best performance.
Error Detection. The results on EgoPER show that our proposed state changes and their counterfactuals can improve performance by combining them all together, verifying the effectiveness of recognizing erroneous labels. In addition, we observe interesting results showing that with only clip-level counterfactuals (ID 3), the model performs slightly worse than its baseline (ID 2). We assume this is because the model overfits short-term procedures and overlook long-term activities procedures and their counterfactual reasoning.
5 Conclusions
In this work, we present a novel procedure-aware video representation learning framework that first incorporates state-change descriptions and state-change counterfactuals in clip-level alignment, enhancing causal reasoning of action transformations. Then, it utilizes video-level counterfactuals that perturb the local actions and create hypothesized scenarios to facilitate the understanding of activity procedures. Our learned representations demonstrate strong effectiveness in terms of procedure awareness and achieve state-of-the-art results on several benchmarks. For future work, a promising direction for further strengthening procedure awareness for video representation learning with state-change counterfactuals is to generate multi-modal counterfactual examples for procedure-aware activities, such as synthetic long-form videos showcasing mistakes and the corresponding corrections.